# The Geometry of Queries: Query-Based Innovations in Retrieval-Augmented Generation

### v1

@@ 1.1 Contributions
    Theoretical Foundation for Query-Space Alignment: We develop a theoretical framework that formalizes the retrieval challenge in RAG systems and elucidates the advantages of query-space alignment. Our analysis underscores the inherent semantic gap between user queries and content embeddings, demonstrating how QB-RAG effectively bridges this gap by operating directly within the query space.
@@
    Survey and Comparative Analysis of Retrieval Strategies: We present a comprehensive survey of prominent retrieval strategies for RAG systems, including naive RAG, QA-RAG, HyDE, and our proposed QB-RAG method. We analyze how each approach addresses the semantic gap between user queries and the knowledge base, comparing their effectiveness through quantitative evaluation on our healthcare dataset. This analysis provides insights into the trade-offs involved in selecting a particular retrieval method for a specific application.
  4. 4.
    Comprehensive Benchmark and Empirical Validation: To evaluate the performance of various retrieval methods, we conduct rigorous experiments on a healthcare question-answering dataset generated from our proprietary healthcare content database. Our evaluation encompasses both retrieval accuracy (using metrics like exact recovery rate and relevance scores) and the quality of the downstream generated answers (assessed using faithfulness, relevancy, and accuracy metrics). This comprehensive benchmark provides valuable insights into the strengths and limitations of each retrieval strategy, showcasing QB-RAG’s superiority on our question-answering dataset. While our evaluation focuses on healthcare question-answering, QB-RAG’s underlying principles and advantages may be broadly applicable to various domains and RAG applications.
The remainder of this paper is structured as follows. Section 2 introduces our healthcare-focused content base and details our methodology for extracting a comprehensive set of questions from these materials. In Section 3, we formalize the RAG objective, highlighting the challenges of retrieval alignment. We then present QB-RAG as a solution, theoretically motivating its effectiveness in bridging the semantic gap between queries and content. Section 4 outlines our benchmark setup, encompassing a range of prominent retrieval methods and evaluation metrics. Finally, Section 5 presents the results of our empirical evaluation, demonstrating QB-RAG’s superiority in retrieval accuracy and downstream answer quality.
2 Application and Dataset
2.1 Description
Digital health programs represent a paradigm shift in managing chronic conditions such as Type 2 diabetes (T2D) and Hypertension (HTN). They can improve health markers associated with chronic conditions by delivering personalized care and support directly to patients (Majithia et al. (2020)). These programs often utilize a variety of digital tools, including mobile apps, messaging, wearable devices, and telehealth services, to facilitate remote monitoring, provide educational content, and offer guidance on lifestyle modifications. The overarching goal is to equip patients with the knowledge and tools necessary to effectively self-manage their condition, enhance their overall quality of life, and potentially reduce the need for frequent traditional healthcare visits.
An integral component of these digital health programs is the messaging platform they incorporate. This platform is not merely a channel for communication; it is an integrated system designed to provide timely, personalized, and interactive support to patients. The effectiveness of these platforms hinges on their ability to engage patients and provide relevant, timely information in response to their queries. However, the increasing demand for immediate healthcare information often strains the capacity of human healthcare providers to deliver timely support through messaging.
To overcome this scalability challenge, digital health platforms are increasingly turning to meticulously curated content repositories designed to provide readily available answers to common patient questions. This study utilizes such a content repository, referred to as “Content Cards,” developed to address frequently asked questions (FAQs) from patients with T2D and HTN participating in our proprietary digital health programs. This dataset, encompassing 630 English-language content cards, covers a comprehensive range of topics pertinent to managing T2D and HTN. This includes general health information, detailed guidance on using platform-specific features and connected devices (e.g., blood glucose monitors), and personalized dietary recommendations. Figure 1 illustrates how these Content Cards are presented within our mobile application, while the excerpt below provides a representative sample of the content:
@@ 2.1 Description
Figure 1: Illustration of Content Cards in Verily’s Onduo mobile application, these cover broad range of topics, from operational health management, nutrition and physical advice, etc.
2.2 Offline Question Generation
The core premise of our work is to improve content retrieval in RAG systems by directly aligning incoming user queries with a pre-computed set of questions derived from our content base. This requires generating a comprehensive set of questions that are answerable by our content, which we achieve through a two-step process. First, we employ a base prompt to generate a preliminary set of questions for each content card. Subsequently, we leverage an LLM-based answerability model (detailed in Section 2.2.2) to filter out irrelevant or nonsensical questions, ensuring the quality and relevance of our question set.
2.2.1 Base prompt
To generate an initial set of questions, we design a prompt (shown in listing 1) with clear instructions for question generation and few-shot examples to illustrate the desired style and format. This prompt incorporates a num_questions parameter to specify the target number of questions per content card. While we set this parameter to 20, the actual number of questions generated by the LLM may vary. Applying this prompt to our content base of 630 cards resulted in over 8,000 potential questions.
1***SYSTEM:***
2You are a Teacher / Professor. Your task is to setup {num_questions} questions for an upcoming quiz/examination. The questions should be both diverse and exhaustive in nature across the document. Restrict the questions to the context information provided.
3
4***INSTRUCTIONS:***
5You are presented with a text authored by healthcare professionals, offering advice and strategies for managing conditions such as diabetes and high blood pressure. Your task is to formulate relevant questions that the text is written to address. Closely follow the example questions for style and structure when formulating your own question for the provided text. Your generated questions should be in first person with details, but only at a high school reading level. Your questions should be answerable from the text, but do not copy the text verbatim. MAKE SURE to generate at least {num_questions} questions. Format the generated questions separated by comma in the following JSON format with "questions" as its key: {{"questions": ["...","...","...","..."]}}
6
7***EXAMPLE OUTPUTS:***
8Sample Text: {example content}
9Generated Questions: {example JSON with list of num_question questions}
10
11--------------------------------------------
12Given the context information and no prior knowledge, generate the relevant questions.
13Text:
14{cc_text}
15Generated Questions:
Listing 1: Base prompt for question generation
2.2.2 Answerability model
After generating questions using an LLM for our content cards, we noticed that our output questions were sometimes not relevant to the content, or were formatted inappropriately. To mitigate this issue and filter out irrelevant questions, we developed an LLM-based answerability model. Specifically, for each content card and generated question pair, we prompt the LLM to assess whether or not the content contains relevant information to answer the question. Leveraging the LLM’s reasoning abilities, we structure the prompt to elicit a step-by-step explanation of the answerability judgment. The LLM is instructed to first identify the key information required to answer the question and then determine if that information is present within the content. The LLM also outputs a “Yes” or “No” label which we use to filter out questions that are deemed irrelevant to the corresponding content card.
Applying this answerability model effectively filters out irrelevant or nonsensical questions generated in the initial phase. This filtering process resulted in a refined set of approximately 4,800 answerable questions derived from our 630 content cards. Each question in this curated set is directly mapped back to its source content card, which serves as the “golden” source for a correct answer.
To validate the effectiveness and reliability of our answerability model, we randomly selected 100 question and content card pairs and evaluated them using both our automated model and three clinical experts. The clinical experts rated the answerability of each pair on a three-point scale: “Content answers completely,” “Content answers partially,” or “Content does not answer.” The results demonstrated strong agreement between our model’s assessments and the clinical expert’s judgments. Specifically, 90% of the pairs received the same answerability rating, with an additional 9% categorized as “partially answerable,” indicating some degree of inherent subjectivity in the evaluation task.
This high level of agreement, coupled with the increasing recognition of LLMs as reliable auto-evaluation tools Lee et al. (2023), supports the validity of our answerability model for filtering irrelevant questions. We provide the prompt used for our answerability model in Listing 2.
1***SYSTEM***
2You are a health coach providing support for members living with diabetes. You have some basic healthcare and nutrition knowledge.
3
4***INSTRUCTIONS***
5Given a pair of user query and a paragraph of content, determine if the content contains relevant information to infer an answer to the query. Think step by step. First provide an explanation, then generate a "Yes" or "No" label. Put the results in a Python dictionary format with keys "Explanation" and "Source relevant".
6
7***EXAMPLES***
8For example, given the following query and content as inputs: {positive and negative examples, with explanations}
9
10--------------------------------------------
11Provide the output for the following query and content:
12Question: {question}
13Content: {content}
Listing 2: Prompt for Answerability Auto-Evaluator
2.2.3 Analysis of Generated Question Similarity

### v2

@@ 4.3.2 Oracle and Approximate QB-RAG
Given the oracle $A^{*}$ or some estimate $\hat{A}$ thereof, we can adapt Algorithm 1 to incorporate the non-sparse nature of $A^{*}$ or $\hat{A}$. Specifically, since multiple content pieces may be associated with a single question, we introduce tie-breaking rules within the algorithm that first prioritize newer content (especially relevant to healthcare care). We then prioritize the content capable of answering a diverse range of questions within $\mathcal{Q}$. This reflects our intuition that such content is likely to be more broadly relevant.
5 Methods
5.1 Offline Question Generation
The core premise of our work is to improve content retrieval in RAG systems by directly aligning incoming user queries with a pre-computed set of questions derived from our content base. This requires generating a comprehensive set of questions that are answerable by our content, which we achieve through a two-step process. All LLM generations utilize Gemini-Pro and text embeddings are generated using textembedding-gecko – both of which are commercially available from Google.
5.1.1 Base prompt
To generate an initial set of questions, we design a prompt (Appendix H.1) with instructions and few-shot examples. This prompt incorporates a num_questions parameter to specify the target number of questions per content card. While we set this parameter to 20, the actual number of questions generated by the LLM may vary. Applying this prompt to our content base of 630 cards resulted in over 8,000 potential questions.
5.1.2 Answerability model
To mitigate issue of irrelevancy of the generated questions to the content, we developed an LLM-based answerability model. For each content card and generated question pair, we prompt the LLM to assess the content’s ability to answer the question. We structure the prompt to elicit a step-by-step explanation of the answerability judgment (Appendix H.2).
This filtering process resulted in a refined set of approximately 4,800 answerable questions derived from our 630 content cards. Each question in this curated set is directly mapped back to its source content card, which serves as the “golden” source for a correct answer.
To validate the effectiveness and reliability of our answerability model, we randomly selected 100 question and content card pairs and evaluated them using both our automated model and three clinical experts. The clinical experts rated the answerability of each pair on a three-point scale: “Content answers completely,” “Content answers partially,” or “Content does not answer.” The results demonstrated strong agreement between our model’s assessments and the clinical expert’s judgments. 90% of the pairs received the same answerability rating, with an additional 9% categorized as “partially answerable,” indicating some degree of inherent subjectivity in the evaluation task. To further validate the answerability model, we additionally confirmed that the semantic similarity among the generated questions above are as expected (Appendix E). This comprehensive analysis, coupled with the increasing recognition of LLMs as auto-evaluation tools Lee et al. (2023), supports the validity of our answerability model for filtering irrelevant questions.
5.2 Benchmarks
[…]
We note that that the methods presented above are not mutually exclusive and can be combined to potentially achieve improved performance, as illustrated by QA-RAG (Kim and Min, 2024). Our experiments focus on evaluating the efficacy of each method in isolation.
5.3 Answer Generation
After retrieving relevant content using the methods described above, we employ an LLM (Gemini-Pro) for answer generation. For each question, the retrieved content is provided as context to the LLM along with the prompt in Appendix H.3 to generate the answer.
5.4 Test Data Generation
We construct two distinct test sets designed to assess the performance of the various retrieval methods under different conditions. Each test set consists of questions answerable by our content base, ensuring relevance to the task. Crucially, no test question appears verbatim within the pre-generated question set used for retrieval. All test questions are generated using Gemini-Pro.
  1. 1.
    Rephrase: To generate this test set, we prompted an LLM to rephrase each of the 4.8k questions in the knowledge base. Then, 500 of those were randomly sampled to yield the first test set. This approach was done to simulate scenarios where the knowledge base contains a comprehensive set of questions. In such scenarios, the intents of new incoming questions are more likely to be represented in the knowledge base, as might occur for mature systems.
  2. 2.
    Out-of-Distribution: To generate the second test set, we prompted an LLM to generate a new question for each of the 630 contents. The LLM was instructed to not generate a question that already exists in the question knowledge base for that content. Then, the 630 newly generated questions were filtered via the answerability model to ensure the new questions were indeed answerable by the corresponding content. After filtering, 305 newly generated questions made up the second test set. While questions in this Out-of-Distribution test set do not inquire the same exact knowledge as the existing questions in the retrieval pool, they are related and can be connected to an existing question in the embedding space. This test set reflects the adversarial, low-coverage scenarios (e.g. systems with a relatively cold start).
6 Metrics
To rigorously assess the performance of our proposed QB-RAG method for healthcare applications, we employ two distinct sets of metrics: those evaluating the quality of content retrieval and those assessing the quality of the generated answers. While QB-RAG’s primary objective is to enhance retrieval, we hypothesize that this improvement will translate to better-generated answers.
6.1 Retrieval Evaluation
  • Auto-evaluator Relevancy Rate: This metric addresses the limitations of relying solely on exact matches by leveraging the answerability model to gauge content relevance. It calculates the percentage of test questions for which at least one retrieved document is deemed relevant by the answerability model. This automated assessment has demonstrated strong correlation with human judgments as as described previously.
6.2 Answer Evaluation
  • Answer Guideline Adherence Rate: The assessment involves a three-step process. Initially, an LLM generates a “golden answer” for each test question using the associated content. Subsequently, another LLM analyzes this golden answer to create a guideline outlining the key elements an accurate response should contain. Finally, a third LLM evaluates the candidate answer against this extracted guideline, assigning a score from 0 to 1 based on the extent to which the answer covers the guideline’s key points. While this guideline-based approach aims to capture the nuances of answer quality, it relies on the accuracy of both the golden answer and the extracted rubric. To provide a more robust and multifaceted evaluation, we introduce additional metrics that directly assess distinct aspects of answer quality below.
  • Answer Faithfulness Rate: This evaluates whether the content supplied to the LLM during generation supports the generated answer. An answer is deemed faithful if any portion of the provided content supports it, regardless of the presence of irrelevant content. This component is crucial to ensure that the answer generated is grounded on our content, a crucial feature for many healthcare applications. We penalize any unwarranted extrapolation or information not grounded in the provided content. Note that declined answers are considered not faithful in our definition.
6.3 Statistical Analysis
To validate the performance differences between models, we conducted statistical significance testing, comparing each model against the NaiveRAG baseline. For binary evaluation metrics, we used a one-sided exact binomial test on the discordant pairs. For the continuous guideline adherence rate, we employed the one-sided Wilcoxon signed-rank test. For all analyses, a p-value of p<0.05 was considered statistically significant.
6.4 Note on Auto-Evaluator Use
For all metrics that require LLMs for annotation (all but exact recovery), we conducted manual annotation to validate the use of auto-evaluators. In section 5.1.2, we separately validated the auto-evaluator relevancy for retrieval on 100 annotations. For answer related metrics, we conducted 2 manual annotations on 50 questions each, ensuring diverse sets, with full results in Appendix B. Altogether the annotation showed 90+agreement between auto-evaluators and human annotation for binary tasks, and high correlation (.84 Spearman rank correlation) on scoring tasks. Inter-rater agreement is comparable, validating the use of auto-evaluators throughout the experiments.
7 Results
7.1 Effect of QB-RAG on Retrieval Efficacy
The exact recovery rate does not present the full picture. When analyzing the Auto-evaluator Relevancy (refer section 6.1), our method still outperforms the baselines when retrieving a single content. As measured by the LLM, the content we retrieve are relevant 68% of the time, whereas traditional methods retrieve relevant content around 40% of the time.
Increasing the number of retrieved documents to 3 further illustrates the efficacy of QB-RAG-8. The exact recovery rate reaches a near-optimal 97%, substantially higher than the baseline approaches. QB-RAG-8 also achieves almost 20% higher relevancy rate compared to the baselines. These findings underscore the robustness of QB-RAG-8 in retrieving both the target document and a set of relevant documents.
Retrieval assessment Rephrase Exact Recovery Rate Auto-evaluator Relevancy Rate
Number of retrieved docs 1 3 1 3
QB-RAG-8 0.89* 0.97* 0.68* 0.76*
QB-RAG-2 0.59* 0.75* 0.53* 0.66*
Naive RAG 0.44 0.61 0.39 0.56
QA-RAG 0.45 0.60 0.41 0.56
HyDE 0.47 0.63 0.41 0.56
Table 1: Retrieval performance (higher is better) of methods on Rephrase, where documents were retrieved given rephrased questions from the content base.* denotes statistical sign.
Furthermore, QB-RAG-8 demonstrates a significant improvement in relevancy. The auto-evaluator relevancy rate shows gains of 8% to 15% over baseline methods, indicating that QB-RAG-8 effectively identifies relevant content even when the query distribution shifts. These improvements in retrieval quality directly benefit downstream answer generation, as more relevant content is likely to lead to more accurate and informative answers.
Retrieval assessment Out-of-Dist Exact Recovery Rate Auto-evaluator Relevancy Rate
Number of retrieved docs 1 3 1 3
QB-RAG-8 0.53 0.72 0.58* 0.75*
QB-RAG-2 0.42 0.57 0.50 0.64
Naive RAG 0.50 0.68 0.50 0.64
QA-RAG 0.51 0.66 0.47 0.60
HyDE 0.52 0.70 0.48 0.65
Table 2: Retrieval performance (higher is better) of methods on Out-of-Distribution, where documents were retrieved given newly generated questions.* denotes statistical sign.
7.2 Effect of QB-RAG on Generated Answer Quality
As shown in Table 3, both QB-RAG-8 and QB-RAG-2 consistently outperform the benchmark methods on all answer quality metrics for the Rephrase test set. Notably, QB-RAG-8 achieves an 84% answer faithfulness rate, significantly surpassing the 62%-68% rates of the baseline methods. This suggests that by accurately retrieving the most relevant content, QB-RAG enables the LLM to generate answers that are well-grounded in the provided information. Additionally, QB-RAG achieves the highest answer guideline adherence rate, indicating its answers effectively address the key elements outlined in the pre-defined guidelines (refer to Section 6.2).
Answer assessment Rephrase Declined Rate Faithfulness Rate Answer Relevancy Rate Guideline Adherence Rate
QB-RAG-8 0.12* 0.84* 0.83* 0.79*
QB-RAG-2 0.21* 0.74* 0.73* 0.73*
Naive RAG 0.30 0.67 0.66 0.68
QA-RAG 0.33 0.62 0.62 0.62
HyDE 0.30 0.68 0.66 0.67
Table 3: Answer quality of methods on Rephrase, where documents were retrieved given rephrased questions from the content base. 3 documents retrieved.* denotes statistical sign.
Answer assessment Out-of-Dist Declined Rate Faithfulness Rate Answer Relevancy Rate Guideline Adherence Rate
QB-RAG-8 0.17* 0.78* 0.77* 0.69
QB-RAG-2 0.29 0.66 0.65 0.63
Naive RAG 0.27 0.72 0.70 0.64
QA-RAG 0.29 0.68 0.67 0.63
HyDE 0.24* 0.74* 0.72* 0.68
Table 4: Answer quality of methods on Out-of-Distribution, where documents were retrieved given newly generated questions. 3 documents retrieved.* denotes statistical sign.
[…]
7.3 Effect of Coverage of Generated Questions
On the Rephrase test set, QB-RAG-2, despite its reduced question set, still surpasses other retrieval methods (table 1). However, the magnitude of improvement is noticeably smaller compared to QB-RAG-8. For instance, QB-RAG-2 shows a 10-15% improvement in exact recovery rate over baselines, whereas QB-RAG-8 nearly doubles the exact recovery rate. This pattern also holds for relevancy rate, indicating that a larger, more comprehensive question set translates to more effective retrieval.
This sensitivity to question coverage is further evident in the quality of the generated answers (Tables 3 and 4). QB-RAG-8 consistently leads to more faithful, relevant, and accurate answers compared to QB-RAG-2, directly reflecting the differences observed in their retrieval performance. These findings highlight that QB-RAG’s success in downstream tasks is fundamentally linked to its ability to construct and leverage a comprehensive and diverse question set that effectively captures the content and semantic nuances of incoming queries within the RAG knowledge base.
7.4 MedQUAD Experiment
We additionally compared the methods with the MedQUAD dataset, which includes pairs of human-authored questions with supporting sources, across diverse healthcare topics (Ben Abacha and Demner-Fushman, 2019). We provide key results in Appendix A, showing up to 20% gain in retrieval accuracy and 10% for final answer quality for select metrics.
8 Discussion
This paper introduces QB-RAG, a framework for enhancing the retrieval phase of RAG systems, with a particular focus on healthcare question-answering. QB-RAG tackles the challenge of semantic misalignment between user queries and knowledge bases by pre-generating a comprehensive set of answerable questions directly from the healthcare content. While the QB-RAG approach shares similarities with prior work on offline query synthesis for information retrieval, its emphasis on generating answerable questions tailored for healthcare, combined with the LLM-based answerability filtering, distinguishes it within the broader landscape of RAG system enhancement techniques. Unlike conventional methods that rely on online LLM calls for query rewriting or enhancement, QB-RAG shifts the computational burden of question generation upfront and offline, thereby reducing latency at inference time. Furthermore, QB-RAG benefits from the efficiency of highly optimized retrieval algorithms, even when handling large question sets, ensuring rapid and practically tractable content retrieval for incoming online queries.
Our experiments demonstrate that QB-RAG outperforms benchmark methods on two distinct test sets: the Rephrase set, which simulates a scenario with comprehensive question generation, and the more challenging Out-of-Distribution set, designed to assess robustness under varying conditions. The results highlight QB-RAG’s ability to improve not only the exact recovery and relevance of retrieved documents but also the quality of downstream answer generation, as evidenced by higher scores on faithfulness, relevancy, and guideline adherence metrics. Crucially, answers generated using QB-RAG are more frequently grounded in the trusted source content, which is important for delivering reliable, up-to-date information to patient queries. Our sensitivity analysis further reveals that the effectiveness of QB-RAG is tied to the coverage and diversity of the generated question set. This suggests that future work could explore more sophisticated prompt engineering, supervised fine-tuning, or RLHF frameworks to enhance the diversity of generated questions. Metrics like BERTScore (Zhang et al., 2019) and Word Mover’s Distance (Kusner et al., 2015) could be used to evaluate this diversity, and an LLM could serve as an auto-evaluator in this regard.
QB-RAG’s strengths extend beyond direct health question-answering, and can be valuable when integrated into EHR systems for telemedecine platforms with appropriate considerations of real-world implications. Several aspects, such as complying with relevant regulations, gauging the need for human supervision in for high-risk queries, and determining ethical deployment paths that track with the system’s readiness for different risk levels (as they relate to misinformation or hallucinations) will be necessary For efficient and accurate functionality of QB-RAG in various healthcare settings.
Despite its strengths and promising results, our work has several limitations that warrant further investigation. First, our evaluation focused on a curated healthcare content base consisting of well-crafted, concise pieces. Assessing QB-RAG’s performance on larger, more diverse datasets, including public datasets or those from other domains and containing less structured or longer-form content, remains an important area for future research. This will establish the generalizability of our findings beyond the specific context of this study. Future research could also expand the system’s capabilities to handle increasingly wider scopes of queries, including unanticipated queries. This could include integrating hybrid approaches with benchmark methods (Appendix F), dynamically updating content base (and therefore question base) from trusted sources, or adding incoming queries incrementally to the deployed system (after matching and validation) to continuously expand the retrieval pool.
Second, the current study relies on LLM-based evaluation metrics. While LLMs have shown promise in assessing answer quality, the real-world deployment of QB-RAG in healthcare settings necessitates thorough expert review for each use case, especially for patient-facing applications. Direct patient interactions demand the highest level of scrutiny to ensure accuracy, safety, and reliability of generated information. Therefore, while automated evaluation serves as a valuable first step, human expert validation remains paramount for clinical deployment.
Third, a broader challenge inherent to all RAG systems, particularly in fast-changing domains like healthcare, is the dependence on the quality and recency of the underlying knowledge base. Medical consensus and best practices are constantly evolving. Thus, content needs to be regularly reviewed and updated. Developing robust mechanisms for ongoing knowledge base maintenance and ensuring alignment with current clinical guidelines is crucial, regardless of the specific retrieval methods employed.
Fourth, while we’ve positioned the offline question generation as an advantage for real-time applications, it’s important to acknowledge that the initial generation and filtering process can be computationally expensive. Future work could explore methods for optimizing this process, such as more efficient filtering mechanisms, more targeted question generation strategies, parallelization (Zhou et al., 2024), active learning, smaller model distillation, or, as indicated above, progressively expanding the retrieval pool leveraging actual queries.
Appendix A MedQUAD Results
We filtered MedQuAD [^1] https://huggingface.co/datasets/lavita/MedQuAD down to 1000 questions of the most prevalent disease areas, then removed 2 duplicates. We used the resulting 998 human derived questions as test queries, and the 114 associated unique websites as our document base. For QB-RAG we generated queries, resulting in 3068 total (recent model versions were more verbose enabling larger knowledge base). We report key results for QB-RAG-27 and QB-RAG-8 by down-sampling.
For fairness across benchmarked methods, we used Gemini-2.0-flash (previous models are unavailable) for all LLM tasks. See table for full results.
  1. 1.
    For 1 document retrieved, QB-RAG-27 and QB-RAG-8 achieve 50-52% exact recovery, while baselines are 31-37%. Document relevancy rate soared from 83-87% to 90-94%!
  2. 2.
    By all metrics, final answer quality improved: Answer relevance (resp. Faithfulness; Guideline adherence) improved from 90-94% (resp 92-95%; 53-56%) to 94-97% (resp 96-98%; 59-63%) using QB-RAG-27 and QB-RAG-8
@@ Appendix A MedQUAD Results
    For 3 documents retrieved, scores are similar across all methods, indicating less obvious benefit from QB-RAG in this specific setting
MedQUAD Human Questions Exact Recovery Rate Auto-evaluator Relevancy Rate Answer Declined Rate Faithfulness Rate Answer Relevancy Rate Guideline Adherence Rate
QB-RAG-27 0.50* 0.94 0.01* 0.98* 0.97* 0.63*
QB-RAG-8 0.52* 0.90 0.03* 0.96* 0.94* 0.59*
NaiveRAG 0.31 0.83 0.06 0.92 0.90 0.53
QA-RAG 0.34* 0.86 0.03* 0.94* 0.91* 0.54
HyDE 0.37* 0.87 0.03* 0.95* 0.94* 0.56
Table 5: Evaluation metrics for different models across Retrieval and Answer quality dimensions on the MedQUAD dataset. Report for 1 document retrieved. * denotes statistical sign.
Altogether, we believe that QB-RAG can improve over SOTA RAG methods in many settings. Yet as seen using 3 documents, this depends on the dataset and hyper-parameters. Nevertheless, combining QB-RAG with existing RAG techniques is promising future work, that would enhance all retrieval by leveraging the best capabilities of each method.
Appendix B Additional Validation Results of Auto-Evaluator
For this annotation task, we first point out that retrieval metrics (section 6.1) are either rule based or already aligned with human labels (see review in section 5.1.2).
Regarding answer evaluation, we constructed 2 separate annotations task to ensure the auto-evaluation (AE) is aligned. We had 2 domain experts annotate the 50 questions for each task (100 total test questions were manually annotated), while blinded to the AE. responses. We separated the tasks to balance out the underlying AE labels given there is class imbalance for each.
  • Task 1: For faithfulness, relevancy and declined-to-answer, we sample a balanced set of 50 auto-evaluated questions (with associated answers and retrieved documents) across our method and different benchmarks (see 5.2). Human evaluators showed 100% agreement with AE on declined to answer (inter-rater agreement 100%), 97% agreement on answer relevancy (inter-rater agreement 98%), 90% agreement on faithfulness (inter-rater agreement 90%).
  • Task 2: For guideline adherence, we sample a set of 50 questions (with associated answer and guideline) with distributed AE scores, across our method and different benchmarks. The 2 domain experts scored the guideline adherence on a 5 point likert scale. Humans evaluators showed .84 Spearman rank correlation with AE (inter-rater correlation .92), and associated .80 Pearson correlation with AE (inter-rater correlation .95).
There were no noticeable differences across methods. Taken together, the above results prove AE is well calibrated (i.e. no more disagreement than within experts) for binary tasks and preserves ranking for quantitative tasks with high correlation. This section boosts are confidence in using AE for assessing retrieval performance and final answer quality.
Appendix C Illustrating Semantic Misalignment on Naive RAG
Here is a failing example for Naive RAG illustrating semantic misalignment between question and document space.
  • $q_{1}:$ What are benefits of soluble and insoluble fiber?
  • $q_{2}:$ What are examples of soluble and insoluble fiber?
  • $c_{1}:$ Soluble fiber helps lower cholesterol and keeps blood glucose from rising quickly. Insoluble fiber helps move food through your body, which can keep your bowels regular. Staying hydrated and maintaining an active lifestyle can also help with digestion and regularity.
  • $c_{2}:$ Oats, beans, lentils, and chickpeas provide soluble fiber. Whole grains, nuts, and vegetables like green beans are good sources of insoluble fiber. Certain fruits like apples and pears also contain both types of fiber.
  • $c_{3}:$ Candy is not a good example of soluble fiber or insoluble fiber.
[…]
Appendix D Content Card Example
@@
Appendix E Analysis of Generated Question Similarity